OpenStudy (anonymous):

Question Part Points Submissions Used Find the area of the region enclosed by the inner loop of the curve. r = 4 + 8sin θ

OpenStudy (anonymous):

equal it to zero first

OpenStudy (anonymous):

0=4+8sin(theta) -4=8sin(theta) -1/2=sin(theta), when is sin theta negative -1/2

OpenStudy (anonymous):

gah i don't have my graphing calculator to check lol

terenzreignz (terenzreignz):

But the question was find the area?

terenzreignz (terenzreignz):

the formula for the area of "under" the curve for polar functions such as this is \[\int\limits_{\alpha}^{\beta}\frac{1}{2}[f(\theta)]^{2}d\]